OpenStudy (anonymous):

solve: 2^2x=8^x-1 for x

OpenStudy (anonymous):

x = 3 do u want explanation???

OpenStudy (anonymous):

2^(2x) = 8^(x-1) 2^(2x) = 2^3^(x-1) 2^(2x) = 2^(3x-3) 2x = 3x - 3 -x = -3 x= 3

OpenStudy (anonymous):

4^x-8^x=-1 lol i was working with this... :D please put appropriate braces...

OpenStudy (anonymous):

thanks guys!

OpenStudy (anonymous):

i m here for u any time but only here ok